OpenStudy (anonymous):

What is the approximate circumference of a semicircle with a diameter of 64 meters? I give medals plz help!!!

OpenStudy (anonymous):

Circumference is 2*pi*R, R is your radius, Diameter is 2x the Radius so 64/2=32. 32 is your radius, from here do 2*pi(3.14)=6.283 and multiply that by 32 so 32*6.283=201.061meters

OpenStudy (maxdrizner):

well half a circle would be 180

OpenStudy (anonymous):

Can I get some more help I have 4 more questions